OpenStudy (ksaimouli):

how to graph sin (pie/2)x

OpenStudy (anonymous):

pi not pie pi=3.14159 pie= tasty treat \[\sin(\frac{\pi}{2})x\] or \[\sin(\frac{\pi}{2}x)\]

OpenStudy (ksaimouli):

the 1st one @completeidiot

OpenStudy (anonymous):

well sin(pi/2) is an actual answer or constant so if you simplify it it will just be a line equation y=cx where c= sin(pi/2)

OpenStudy (ksaimouli):

ohk thx i have another question